# TableLoRA: Low-rank Adaptation on Table Structure Understanding for Large Language Models

### v1

@@ Abstract
TableLoRA: Low-rank Adaptation on Table Structure Understanding for Large Language Models
Xinyi He1thanks: The contributions by Xinyi He and Yihao Liu have been conducted and completed during their internships at Microsoft.   Yihao Liu2*   Mengyu Zhou3thanks: Corresponding author.   Yeye He 3 Haoyu Dong 3   Shi Han 3   Zejian Yuan1   Dongmei Zhang3 1 Xi’an Jiaotong University 2 Peking University 3 Microsoft Research hxyhxy@stu.xjtu.edu.cn, haoeliu@stu.pku.edu.cn, yuan.ze.jian@xjtu.edu.cn, {mezho, yeyehe, haoyu.dong, shihan, dongmeiz}@microsoft.com
@@ 1 Introduction
To promote the recognition and understanding of table structures during tabular tasks and enhance the ability of Large Language Models (LLMs) to process tables within PEFT, we propose TableLoRA [^1] The code will be publicly available upon acceptance, a module compatible with the PEFT framework for LLMs. Existing tabular LLMs attempt to learn table structure relationships using attention mechanisms through additional training. In contrast, we directly inform the model of these relationships through our design, which consists of two key components:
@@ 2.4 Parameter-Efficient Fine-Tuning
Model HiTab WikiTQ FeTaQA TabFact
acc $\Delta$ acc $\Delta$ bleu $\Delta$ acc $\Delta$
Llama 2 Full Finetune 48.61 -0.33 46.20 5.74 32.10 4.10 52.29 -25.77
Lora 43.00 -5.94 38.76 -1.70 25.13 -2.87 76.93 -1.13
TableLoRA 48.94 0.00 40.46 0.00 28.00 0.00 78.05 0.00
Llama 3 Full Finetune 61.62 3.05 56.84 3.39 34.86 4.62 84.18 0.17
Lora 57.06 -1.50 51.98 -1.47 29.09 -1.14 83.49 -0.52
TableLoRA 58.56 0.00 53.45 0.00 30.23 0.00 84.01 0.00
DeepSeek Full Finetune 54.92 7.99 47.01 6.59 32.15 4.87 79.44 2.40
Lora 43.25 -3.69 37.34 -3.08 26.68 -0.61 75.20 -1.85
TableLoRA 46.94 0.00 40.42 0.00 27.29 0.00 77.05 0.00
@@ 4 Experiment
All experiments were run on Linux machines with 4 NVIDIA Tesla A100 80G memory GPUs.
@@ 4.1 Experiment Setup
We conducted experiments on three open-source LLMs: Llama 2 Touvron et al. (2023), Llama 3, and DeepSeek DeepSeek-AI (2024). Details are provided in §A. We use the official metrics for each dataset, as shown in the header of Table 1.
@@
4.2 Main Results
4.3 Control Experiments
4.3.1 Table Representation learning Methods
To further validate the superiority of the TableLoRA structure, we summarized common table representation learning methods in §2.2 and designed controlled experiments based on those designs. The results are shown in Table 2. Referring to the table pre-training methods before the era of LLMs Dong et al. (2022), common table representation learning methods that can be applied to LLMs include: describing through strings, enhancing structural information through positional embedding, and enhancing structural information through attention masks. Therefore, we designed the following variant experiments, and more details are shown in §D.
@@ 5 Conclusion
Appendix A Models Used
@@ Appendix E Details of Further Analysis
E.1 Performance of 2D LoRA at Different Layers

### v2

@@ Abstract
TableLoRA: Low-rank Adaptation on Table Structure Understanding
for Large Language Models
Xinyi He1thanks: The contributions by Xinyi He and Yihao Liu have been conducted and completed during their internships at Microsoft.   Yihao Liu2*   Mengyu Zhou3thanks: Corresponding author.   Yeye He 3 Haoyu Dong 3   Shi Han 3   Zejian Yuan1   Dongmei Zhang3 1 State Key Laboratory of Human-Machine Hybrid Augmented Intelligence, Institute of Artificial Intelligence and Robotics, Xi’an Jiaotong University 2 Peking University 3 Microsoft hxyhxy@stu.xjtu.edu.cn, haoeliu@stu.pku.edu.cn, yuan.ze.jian@xjtu.edu.cn, {mezho, yeyehe, haoyu.dong, shihan, dongmeiz}@microsoft.com
@@ 1 Introduction
To promote the recognition and understanding of table structures during tabular tasks and enhance the ability of Large Language Models (LLMs) to process tables within PEFT, we propose TableLoRA [^1] The code will be open-sourced on https://github.com/microsoft/TableLoRA., a module compatible with the PEFT framework for LLMs. Existing tabular LLMs attempt to learn table structure relationships using attention mechanisms through additional training. In contrast, we directly inform the model of these relationships through our design, which consists of two key components:
@@ 3.1 Special Tokens Encoder
Model HiTab WikiTQ FeTaQA TabFact
acc $\Delta$ acc $\Delta$ bleu $\Delta$ acc $\Delta$
Llama 2 Full Finetune 48.61 -0.33 46.20 5.74 32.10 4.10 52.29 -25.77
Lora 43.00 -5.94 38.76 -1.70 25.13 -2.87 76.93 -1.13
TableLoRA 48.94 0.00 40.46 0.00 28.00 0.00 78.05 0.00
Llama 3 Full Finetune 61.62 3.05 56.84 3.39 34.86 4.62 84.18 0.17
Lora 57.06 -1.50 51.98 -1.47 29.09 -1.14 83.49 -0.52
TableLoRA 58.56 0.00 53.45 0.00 30.23 0.00 84.01 0.00
DeepSeek Full Finetune 54.92 7.99 47.01 6.59 32.15 4.87 79.44 2.40
Lora 43.25 -3.69 37.34 -3.08 26.68 -0.61 75.20 -1.85
TableLoRA 46.94 0.00 40.42 0.00 27.29 0.00 77.05 0.00
Large Models GPT-4o 55.05 - 58.38 - 14.34 - 44.01 -
Claude-3.7 56.88 - 66.90 - 13.58 - 80.14 -
Llama 3.3-70B 22.79 - 41.30 - 7.66 - 15.99 -
@@ 4.1 Experiment Setup
We conducted experiments on three open-source LLMs: Llama 2 Touvron et al. (2023), Llama 3, and DeepSeek DeepSeek-AI (2024). In addition, we include several larger models as baselines for comparison. Due to their closed-source nature or computational constraints, we perform inference only on these models without fine-tuning. Details are provided in §A.2. We use the official metrics for each dataset, as shown in the header of Table 1.
All experiments were run on Linux machines with 4 NVIDIA Tesla A100 80G memory GPUs. The LLaMA Factory framework served as the foundation, which we extensively customized to incorporate TableLoRA-related techniques and methods. In the main experiments, LoRA and 2D LoRA share the same rank (8). PEFT is applied to the $k\_proj$ and $v\_proj$ layers. For more training details, see §B.
4.2 Main Results
Compared with large models, on WikiTQ, Claude 3.7 and GPT-4o generally outperform TableLoRA and related baselines, indicating that tuning-based approaches still require improvement compared to reasoning-focused LLMs, and may need enhancement through incorporation of reinforcement learning for further progress. In contrast, on FeTaQA, TableLoRA and its baselines consistently surpass reasoning LLMs, potentially because the tuning process enables better acquisition of FeTaQA’s answer sentence structures, thereby achieving higher BLEU scores.
4.3 Control Experiments
4.3.1 Table Representation learning Methods
To further validate the superiority of the TableLoRA structure, we summarized common table representation learning methods in §2.2 and designed controlled experiments based on those designs. The results are shown in Table 2. Referring to the table pre-training methods before the era of LLMs Dong et al. (2022), common table representation learning methods that can be applied to LLMs include: describing through strings, enhancing structural information through positional embedding, and enhancing structural information through attention masks. Therefore, we designed the following variant experiments, and more details are shown in §5.
@@ 4.4.2 Tables of Varying Complexity
When both the top and left header depths are 1, a slight drop is observed. (1) The HiTab dataset is imbalanced, containing only 35 samples with one-hierarchy headers. This small sample size limits the reliability of statistical conclusions. Case studies of the three misclassified examples indicate that the errors were caused by perturbations. (2) To further assess the effectiveness of TableLoRA on flat tables—including those with one-hierarchy headers—we conducted experiments on additional flat-structured datasets (e.g., WikiTQ, FeTaQA, TabFact, see Table 1), where TableLoRA consistently achieves stable improvements.
@@ 4.4.3 Queries with Different Aggregations
4.4.5 TableLoRA with Different Rank
Figure 6 illustrates the performance of TableLoRA and LoRA under different ranks, demonstrating that TableLoRA consistently improves with varying ranks, indicating the scalability of the method.
@@ 5 Conclusion
Figure 6: Performance of TableLoRA and LoRA with Different Ranks on HitTab. Accuracy is in %.
@@
Acknowledgments
We thank all anonymous reviewers for their valuable comments. Xinyi He and Zejian Yuan were supported in part by the National Key R&D Program of China (2023YFB4704900) and NSFC (61976170, 62088102).
Appendix A Experiment Configuration Details
A.1 Datasets Selection
We selected four datasets (HiTab, WikiTQ, FeTaQA, TabFact) because they are classic, reliable, and cover diverse table structures and tasks. Table 4 summarizes their characteristics. Varying instruction systematically test the model’s ability to locate and retrieve information from tables at different difficulty levels, while diverse output rigorously assess its capacity to generate responses with varying linguistic or structural demands.
Table 4: Dataset Selection.
Dataset Table Instruct Output
HiTab hierarchical query words
WikiTQ flat query words
FeTaQA flat query sentences
TabFact flat fact/statement bool
A.2 Models Used
@@ Appendix D Details of Control Experiment
Table 5: Ablation Study on each model for each dataset. All metric numbers are in %.
Model HiTab WikiTQ FeTaQA TabFact
Llama 2 TableLoRA 48.94 40.46 28.00 78.05
w/o Special Token Encoder 47.19 39.73 27.72 78.00
w/o 2D LoRA 44.13 39.60 25.91 77.37
LoRA 43.00 38.76 25.13 76.93
Llama 3 TableLoRA 58.56 53.45 30.23 84.01
w/o Special Token Encoder 58.43 53.52 29.81 83.88
w/o 2D LoRA 57.25 52.75 30.19 83.23
LoRA 57.06 51.98 29.09 83.49
DeepSeek TableLoRA 46.94 40.42 27.29 77.05
w/o Special Token Encoder 45.63 38.87 27.50 76.78
w/o 2D LoRA 44.44 37.71 26.92 75.62
LoRA 43.25 37.34 26.68 75.20
@@ Appendix E Details of Further Analysis
E.1 Ablation Study
The ablation results for each dataset are shown in Table 5.
E.2 Performance of 2D LoRA at Different Layers